TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-08 23:31:00
subject: 2\16 ISS On Orbit Status 16-02-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

ISS On-Orbit Status 2/16/03

All ISS systems continue to function nominally, except as noted 
previously or below. A quiet Sunday with lots of rest time for the 
crew, including an expanded Sunday-type cartoon section uplink. Ahead: 
Week 13.

FE-1 Nikolai Budarin completed the weekly routine tasks of collecting 
SP toilet flush counter and SVO water supply readings for calldown, as 
well as the periodic inspection of the "Elektron" oxygen generator's 
VM gas/liquid system for the air bubble that usually lingers after an 
in-flight maintenance.

CDR Ken Bowersox performed the daily routine maintenance of the SOSh 
life support system, incl. ASU toilet subsystem, and FE-2/SO Don 
Pettit prepared the IMS inventory database for automated file 
import/export.

Pettit also took the daily ppCO2 (carbon dioxide partial pressure) 
readings in Lab and SM with the CDMK (CO2 monitoring kit), after 
breakfast (1:40am EST) and again before dinner (3:00 pm).

All crewmembers performed their daily physical exercise program on 
TVIS (treadmill with vibration isolation and stabilization), RED 
(resistive exercise device), CEVIS (cycle ergometer with vibration 
isolation), and VELO (Russian cycle ergometer) with load trainer.

The crewmembers' weekly PFCs (private family conferences) were 
scheduled today for Budarin on Russian VHF early in the morning, and 
for Bowersox and Pettit on S-band later in the day.

Sox and Don also had their regular periodic PPCs (private 
psychological conference).

Science Update (Expedition Six -- 11th):

MSG (Microgravity Science Glovebox): ESA and Astrium suspect the MSG 
failure to be due to one or several sensors in the electronics box. 
For testing the sensors, MSG will have to be powered up, but only when 
it is fully understood what damaged the PDC (power distribution 
controller), to prevent damage repetition. NASA and ESA specialists 
are currently looking into design details, with earliest possible 
troubleshooting not before late next week.

Science Downmass: Plans are being made for returning Expedition 6 
science on Soyuz. Requested downmass consists of Renal (kidney) 
collection samples and a subset of the 29 processed ZCG (Zeolite 
crystal growth) autoclaves. Five individually wrapped autoclaves, with 
motor removed, have been specified, in priority order for descent, but 
the Soyuz 5S descent manifest is still very much in flux.

Extra-Vehicular Activity Radiation Monitors (EVARM): Continuing.

GASMAP/Pulmonary Function in Flight (PuFF): The next session is in 
late February.

Renal (Kidney) Stone Experiment: Completed for Increment 6. Bowersox 
and Pettit continue to take pills.

Human Research Facility/Workstation (HRF WS): Continuing.

Foot/Ground Reaction Forces During Space Flight (FOOT): Foot team has
received all data form the last run. The PI (Principal Investigator) 
is reviewing the data and will have words for the crew shortly.

Pore Formation and Mobility Investigation (PFMI): On hold pending MSG 
PDC (power distribution controller) and ESEM3 (exchangeable standard 
electronic module 3) replacement. Will remain in MSG until further 
notice.

Space Acceleration Measurement System (SAMS): SAMS continues to 
collect acceleration data for vibratory characterization of the 
microgravity environment, including the recent reboost activities.

Microgravity Acceleration Measurement System (MAMS): MAMS continues
measurement of microgravity environment in the quasi-steady regime for
general characterization.

Protein Crystal Growth-Single Locker Thermal Enclosure System 
(PCG-STES):  Temperatures are nominal. PCG-STES is currently in 
heating mode.

Investigating the Structure of Paramagnetic Aggregates from Colloidal
Emulsions (InSPACE): On hold until the MSG (microgravity science 
glovebox) is operational.

Materials ISS Experiment (MISSE): In progress. Deployed outside. 
Nominal and collecting data.

Zeolite Crystal Growth (ZCG): ZCG has finished science operations for 
Inc 6.

EarthKAM (EK): All activities complete for Increment 6.

Crew Earth Observations (CEO): Until further notice, the crew is to 
use the digital camera systems only for CEO targets. Their recent 
image of reefs and volcanoes sequence in French Polynesia and a 
striking night view of London have been selected for publication in 
Earth Observatory. Now that light is returning to the Northern 
Hemisphere, the extent and rapid seasonal change of winter's snow and 
ice patterns will offer good targets for Earth photography by the ISS 
crew. But while in XPOP attitude, use of the Science Window in the Lab 
is constrained by Flight Rule.

Today's CEO targets, accounting for XPOP window constraints, were 
Mexico City, Mexico (DYNAMIC EVENT TARGET: Reports early Friday 
morning indicated that Mexico's Popocatepetl volcano had just erupted 
with a 3-mile high column of smoke. ISS track was over Mexico City 
itself, but the volcano is south of the city. Crew was to look right 
of track to document this event), Tigris-Euphrates, Turkey (with 
improving weather at the time of this pass, the crew should have had 
an excellent view right of track of the Tigris-Euphrates Rivers in sun 
glint, especially over Iraq), Albuquerque, New Mexico (although 
Albuquerque is nearly at nadir, crew was advised to look left of 
track, northward toward the Sangre de Cristo and the Front Range of 
the Rockies to document snow pack conditions), Labrador Ice Floes 
(massive ice floes are swept down the coast of Labrador during late 
winter and typically achieve their maximum extent in February. For 
best views, looking obliquely left of track where clouds and weather 
permit, and trying to include landmarks where possible), Sierra Snow 
Pack (from the crew's track position over northern California, they 
had an excellent oblique view down the spine of the High Sierras to 
their right, for documenting conditions of the snow pack), and James 
Bay Ice Pack (this is a frequently encountered target for ISS orbits. 
Crew was asked to begin documenting the state of pack ice in James Bay 
to the left of track, especially near Akimiski Island).

CEO images can be viewed at the website http://eol.jsc.nasa.gov

ISS Orbit (as of this morning, 4:19am EST [= epoch]):
Mean altitude -- 392.1 km
Apogee -- 400.5 km
Perigee -- 383.8 km
Period -- 92.40 min.
Inclination (to Equator) -- 51.63 deg
Eccentricity -- 0.001236
Orbits per 24-hr. day -- 15.58
Mean altitude loss in last 24 hours -- 190 m
Revolutions since FGB/Zarya launch (Nov. '98) -- 24215

For more on ISS orbit and worldwide naked-eye visibility dates/times, 
see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

 - End of File -
================

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.